OpenStudy (gg):

how many zeros polynomial P(z)= 3z^6 +z^5+ z^4+ 2z^3 + 10z^2+ z+1 has in area 1<|z|<2?

OpenStudy (anonymous):

lalay is typing a reply..

OpenStudy (anonymous):

tell me how to begin this problem :P

OpenStudy (gg):

I have to see step by step of solving, I can't just write a number of zeros on exam :D

OpenStudy (lalaly):

lol wait im thinkin

OpenStudy (anonymous):

lalaly is now thinking.. please wait

OpenStudy (gg):

hahahah :D ok :D you are lalaly's lawer? :D

OpenStudy (lalaly):

i tried but i got the answer wrong...im sorryyy

OpenStudy (anonymous):

how to start.. tell me how to start :'(

OpenStudy (gg):

I don't know :D If I know how to start, I;ll probably know how to finish :D I posted another one problem :) are u maybe good in probabillity?

OpenStudy (zarkon):

there are 4, but I used my calculator to get that answer

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

by the looks of this question, although there is no indication given, i think this is a polynomial of a complex variable and 1<|z|<2 means the annulus i think

OpenStudy (zarkon):

all 6 roots are complex

OpenStudy (anonymous):

oh ok. isn't there a formula that counts the number of zeros minus the number of poles?

OpenStudy (anonymous):

and since this obviously has no poles it is just he number of zeros. you integrate p'(x)/p(x) maybe

OpenStudy (zarkon):

I am unfamiliar with that

OpenStudy (anonymous):

i think so it is like residue theorem. divide by 2 pi i too.

OpenStudy (zarkon):

rouche's theorem

OpenStudy (anonymous):

problem is i forget how to integrate this thing around the annulus!

OpenStudy (anonymous):

maybe you are not supposed to integrate. maybe you are supposed to bound then?

OpenStudy (anonymous):

god i have forgotten everything. lucky i remember my phone number

OpenStudy (anonymous):

i am going to try this and see if it works. show that all 6 zeros are inside |z|<2 and then show 2 are inside |z|<1 but i may not be able to my guess is that is the method though. that or integrating \[\frac{1}{2\pi i}\int_{1<|z|<2} \frac{p'(z)}{p(z)}\]

OpenStudy (anonymous):

couldn't make first method work. help!!

OpenStudy (zarkon):

ok... if |g(z)|<|f(z)| for all z on gamma(gamma a simple loop), then f and f+g have the same number of roots inside gamma let \[f(z)=3z^6\] and \[g(z)=z^5+ z^4+ 2z^3 + 10z^2+ z+1\] for \[|Z|=2\] \[|f(2)|=3(2)^6=192\] and \[|g(2)|=2^5+ 2^4+ 2\cdot 2^3 + 10\cdot 2^2+ 2+1=106\] thus \[p(z)\]has the same number of roots as \[3z^6\] which is 6 not let \[f(z)=10z^2\] and \[g(z)=3z^6+z^5+ z^4+ 2z^3 + z+1\] then for |Z|=1 |f(1)|=10 and |g(1)|=3+1+1+2+1+1=9 \[10z^2\] has 2 roots in|Z|<1 thus P(Z) also has 2 roots hence P(z) has 6-2=4 roots in 1<|Z|<2

OpenStudy (anonymous):

oh very nice! i got the first part with the 2 (by copying right off of wiki) but didn't get the second part with the 10z^2.

OpenStudy (zarkon):

well I learned something new ;)

OpenStudy (anonymous):

i guess i could have could have thought somewhat harder about the 10z^2 part since i wanted to show that there were TWO zeros inside |z|<1 ! nice work

OpenStudy (anonymous):

i learned something new as well. happens every once in a while

OpenStudy (zarkon):

:)

OpenStudy (anonymous):

hope Gg comes back and they can learn something as well

OpenStudy (zarkon):

yes

OpenStudy (gg):

thank you very much :) but how do I know what tochoose for f(z) and what for g(z)?